A Bibliography of Publications by, and about, John Todd

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 November 2023 Version 1.03

Abstract

This bibliography records publications of John Todd.

Title word cross-reference

(1 + x)^{−1} [Tod86]. $10 [Tod70a]. $12.50 [Hou63]. 16 × 16 [Tod53]. $6 [Fla64]. $9.75 [Tod56a]. [−1, 1] [CT83b]. [0, 1] [Tod86]. cos z = az + c [CT36]. exp z^{2} ∫_{0}^{z} exp(−t^{2}) dt [KT59a]. J_{3} [Koj78]. ≤ n − 2 [CT83b]. log n [KR79]. QR [KT76]. ℘(z|e_{1}, e_{2}, e_{3}) [Tod90b]. x^{n} − nσx^{n−1} [CT83b].

13-15 [Cra87]. 15th [Jun64]. 16-19 [MTTT63a, MTTT63b]. 1954 [HT91]. 1962 [MTTT63a, MTTT63b]. 1969 [Tod70a]. 1981 [BW83]. 1987 [Cra87]. 1993 [BCEP94]. 1996 [Coh97].

20th [WH97]. 20th-Century [WH97].

3rd [BW83].

67 [Tod70c].

A.M.S [Jun64]. Abh [Tod70c]. absolutely [Tod35, Tod36]. Academic [Fla64, Tod70a]. Accounting [Com36, Tod74]. ACM [Cra87]. ADI [GT67, GT67, Tod67c]. ADI-Parameters [Tod67c]. Admiralty [ST46, TS47]. Advanced [BT50, ET46, NST52]. Advances [Tod63a]. AGM [TBDS92a, TBDS92b]. Akad [Tod70c]. Alfred [Tod50a]. Algebra [Tod77, TT47, Ran80, Row79]. Algebraic [Tod58a, FTT55a, TT40a]. Algebraic-Hyperbolic [Tod58a]. Algebras [Tod50a]. Allied [Tod59b]. Almanac [Tod59b]. Alt [Tod63a]. Alwin [Tod93]. American [MTTT63a, MTTT63b]. analogs [FTT55b]. Analysis [HT91, Ran80, Tod62d, Tod78a, Mad12, Tod50c, Tod54f, Tod55a, Tod55c, Tod57a, Tod57b, Tod59d, Tod62a, Tod62c, Tod75c, Hou63, Tod55a, Tod57a, Tod57b, Col64, Olv63, Pra63, Tod56a, Too63, Tod58b, Tod67a]. Analytic [Tod61b]. ancient [Tod87]. angewandte [Tod52]. Answer [TS85]. Applications [Com36, TTT57, TT57a, TT12a, Tod84, Tod80a]. Applied [Jun64, MTTT63a, MTTT63b, Tod56a, Tod58b]. approach [Tod56b]. approximation [CT83b, Tod86]. Approximations [Tod70a]. April [BW83, Coh97, MTTT63a, MTTT63b]. arc [Tod49b]. Arctangents [Tod51]. arguments [Tod54d]. Arithmetic [Ham65b, Ham65a, Ham65c, Jun64, MTTT63a, MTTT63b, Tod37]. Army [Tod70a]. aspects [TT47]. Asymptotic [QZL^{+}10]. Atlantic [MTTT63a, MTTT63b]. Automatic [NT62, Tod50e].

B [Ham65b, Ham65a, Ham65c, Jun64, Tod67a, Tod67b, TBDS92a, TBDS92b]. Basel [Tod52]. Basic [Tod77, Tod78a, Ran80, Row79]. Bayer [Tod70c]. Begründung [Tod55a, Tod57a, Tod57b]. Behavior [CT83a]. Bernice [Tod54g]. Beschäftigung [Tod55a, Tod57a, Tod57b]. Bessel [KT60]. best [CT83b, Tod86]. between [MTTT63c]. Bibliography [OTT49]. bilinear [TT56a]. Bing [Tod58a]. Birkhäuser [Tod52]. Bjarni [Tod50a]. Boas [Tod61b]. Bogel [Tod59a]. Book [Col64, Fla64, Ham65b, Ham65a, Ham65c, Hou63, Jun64, Mil65, Pra63, Ran80, Row79, Tod38, Tod40, Tod50a, Tod52, Tod56a, Tod58b, Tod58a, Tod59a, Tod59c, Tod59b, Tod61a, Tod61b, Tod63a, Tod64, Tod67a, Tod70a, Tod71, TBDS92a, TBDS92b, Too63, BBB00].

Booklet [Tod59b]. Brown [Tod54g]. Buck [Tod61b]. Bureau [Tod75c, Tod87, Tod90a].

C [Ham65b, Ham65a, Ham65c, Jun64, Tod61b, TBDS92a, TBDS92b]. Cardinal [Tod50a]. Carlo [Tod54e]. Carlson [TBDS92a, TBDS92b, Tod88b]. Carolina [BCEP94]. Cauchy [ST51]. Cauer [Tod70c, Tod67b]. Cauer-Parameter [Tod70c]. Centenary [BCEP94]. Center [Tod70a]. Century [WH97]. certain [Tod49a, Tod50b, Tod54b, Tod58c, Tod60]. characterisation [TT40a]. characteristic [TT51]. Chebyshev [Tod67b]. Chicago [MTTT63a, MTTT63b]. Chih [Tod58a]. Chih-Bing [Tod58a]. Cholesky [TT06]. City [MTTT63a, MTTT63b]. Classical [Tod62a]. cm [Hou63]. cm. [Tod59c]. Co [Hou63]. Comment [Tod54g, TS85]. Comments [Tod93]. Commuting [TT56a]. Companion [Tod59b]. complex [Tod54d]. Comprehensive [Tod70b]. Computation [Cra87, Tod55b, Tod65b, Tod87, Tod90a]. Computational [Tod61c, TW55]. Computations [New62, TT60]. computers [NT62, Tod63a]. Computing [Ham65b, Ham65a, Ham65c, Jun64, MTTT63c, Nas90, ST46, TS47, Tod50e, MTTT63a, MTTT63b]. concerning [Tod61c]. Condition [Tod54c, Tod68, Tod49a, Tod50b, Tod54b, Tod58c, Tod60]. Conference [BW83, BCEP94, Cra87]. conformal [Tod55b, TW55]. Constants [Tod75a, Tod75b, Tod00, LT72, LT73]. construction [Tod59c]. Constructive [Fla64, Mil65, Tod63b, Tod62b, Tod59a]. continuous [Tod35, Tod36]. Contributions [Tau54]. Convergence [ST51, GT67, GT67]. converse [Yin94]. Converses [Alz91]. Cornelius [BCEP94, Tod58b, Tod94b]. Corrigendum [Tod75a]. Covering [TT48a, TT48b, TT49]. cubic [Tau87]. Curtiss [Tod56a, Tod80b]. curves [Tau87].

Debate [Tod93]. December [BCEP94]. Deductions [ST46]. Degenerating [CT83a]. degree [CT83b]. density [CT49]. determinantal [FT55, Tau87]. Determinants [TTT57, TT40b, Tod54g, TT52a, TT52b, TT57a, TT57b, TT12a, TT12b]. Determination [Tau54]. development [Mad12]. Dies [WH97]. Differential [Tod50e, Tod64, Tod50c, Tod54e, Tod56b]. Differenzenrechnung [Tod61a]. digital [Tod65b]. digitale [Tod52]. dimensions [KT76]. direct [Tod56b]. Discrete [FTT55b, TT60]. Discrimination [QZL^{+}10]. Discussions [Tod50e]. disks [Tod65a]. Distance [QZL^{+}10]. distances [Tod40]. d’un [Tod40].

E. [Tod88a]. E3219 [RKB^{+}87]. E3224 [RKB^{+}87, RT87, RTDH90]. E775 [BT48]. E852 [DFT49]. early [Tod90a]. Ed [Tod56a, Tod70a]. Editor [Hou63, Tod94a]. Eigenvalues [Tau54, Tod65a]. elektronische [Tod52]. Elementary [BT48, DFT49, RKB^{+}87, RT87, RTDH90]. elements [TT51]. Elliptic [CT83a, Tod80a]. Emphasis [Tod70a]. engaging [Tod55a, Tod57a, Tod57b]. England [Tod59c]. ensembles [Tod40]. equation [TW55]. Equations [TTT57, Tau54, Tod50e, Tod54g, TT52a, TT52b, TT57a, TT57b, TT12a, TT12b, Tod50c, Tod54e, Tod56b, Tod60, Tod64]. Ernst [Tod70c]. erratum [Tod70c]. error [KT58, KT59b, Tod65b]. espace [Tod40]. euclidien [Tod40]. Evaluation [KR79, Tod54d, NT58]. Expansions [Tod61b]. Experience [ST46]. Experimental [Ham65c, Jun64, MTTT63a, MTTT63b, Ham65b, Ham65a]. Experiments [Tod53, Tod54e, Tod55b, TW55]. exponential [Tod54d]. extrema [Tod50d].

F [Tod54a]. factorization [TT06]. Fan [Alz91, Lun94, Yin94]. fifteenth [MTTT63a, MTTT63b]. Finite [Tod54c, TT40c, TT41b]. first [CT83b]. Formulae [Tod71]. Fox [Tod59c, Tod64]. fr. [Tod52]. Franz [Tod63a]. function [KT58, KT59b, KT59a, Tod59a]. Functions [CT83a, Tod63b, Tod70a, KT60, Tod35, Tod36, Tod50d, Tod62b, Tod80a, Fla64, Mil65, Tod61b].

G. [Tod94a]. Gelfond [Tod61a]. General [BW83, BW83]. generalization [Lun94]. Generation [TT54, TT56b]. German [Tod55a, Tod57a, Tod57b]. Gerschgorin [Tod65a]. Gershgorin [TW55]. Ghizzetti [Tod71]. Glowatzki [Tod70c]. Government [ST46]. groups [TT41a, TT48a, TT48b, TT49].

H [Ham65b, Ham65a, Ham65c, Jun64,

Tod52, Tod56a, Tod67a, Tod94a]. Hamilton [Tod80b]. Hardy [Tod94a, TV99]. Heft [Tod70c]. held [MTTT63a, MTTT63b]. Her [Tod59c, Tod59b]. High [Ham65b, Ham65a, Ham65c, Jun64, MTTT63c, MTTT63a, MTTT63b]. High-Speed [MTTT63c]. Hilbert [Tod54c, Tod61c]. Hill [Hou63, Tod56a]. Historical [Tau87]. History [Cra87, Nas90, Tod87, Tod90a]. HSNC’87 [Cra87]. Hyperbolic [Tod58a].

II [TT52b, Tod36, Tod54b, TW55, Tod88b]. III [Tod58c]. Illinois [MTTT63a, MTTT63b]. Illus [Fla64, Tod70a]. INA [HT91]. Industry [ST46]. Inequalities [BW83, Tod67b, Alz91, FTT55b, Lun94]. inequality [FTT55a, FT55, TV99, Yin94, TBDS92a, TBDS92b]. Infinite [TT42]. Institut [Tod52]. Institute [HT91]. Instruction [ET46]. integers [CT49]. integral [TT51, Tod54d, TW55, Tod60, Tod38]. Integrals [Tod58a, ST51, Tod80a]. Interactions [MTTT63c]. International [BW83, BCEP94]. Interpolation [Com36, Tod59b]. Interview [Coh97]. Introduction [Fla64, Tod63b, Mil65]. Inverse [Com36]. Inversion [TT41a, Tod53, NT58]. Isaacson [Tod67a]. isolated [Tod65a]. isoperimetric [FTT55a].

J [Tod70a]. J. [Alz91]. Jack [OR07]. Jersey [Cra87, MTTT63a, MTTT63b]. John [Ano97, Col64, Fla64, Ham65b, Ham65a, Ham65c, Jun64, Mil65, Pra63, Ran80, Too63, Alb07, Bau07, Coh97, Hou63, Mad12, Olv63, Tau87, Tod74, Tod80b, WH97]. Jónsson [Tod50a]. Jordan [Tod67b]. Jr [Tod61b].

Kac [TT91]. Keller [Tod67a]. Kl [Tod70c]. Ky [Alz91].

L [Tod63a, Tod64]. L. [Tod54a, Tod59c]. Lab [Tod59c]. Laguerre [Tod50d]. Lanczos [BCEP94, Tod58b, Tod94b]. large [CT36, Tod54d]. legacy [Tod84, Tod88a]. Lemniscate [Tod75a, Tod75b, Tod00]. Lichtenstein [TW55]. limits [Tod78b, Tod88b]. Linear [Tau54]. Ling [Tod58a]. loci [Tau87]. Logarithmic [TBDS92a, TBDS92b]. London [Tod59c]. look [TT91].

M [TBDS92a, TBDS92b]. Machine [Com36, Tod74]. Machinery [Tod50e]. Madison [Tod70a]. Majesty [Tod59c, Tod59b]. many [Tod78b, Tod88b]. mapping [TW55]. maps [Tod55b]. March [Coh97]. Mark [TT91]. Math [Tod59c, Tod70c]. Math.-Natur [Tod70c]. Mathematical [MTTT63a, MTTT63c, MTTT63b, Tod59c]. Mathematics [ET46, Fla64, JB57, JB12, Jun64, Ran80, Row79, ST46, Tod56a, Tod70a, Tod77, Tod78a, WH97, Alb07, MTTT63a, MTTT63b, Ham65c, Ham65b, Ham65a, Jun64]. Mathematik [Tod52]. Matrices [TTT57, TT40c, TT41b, Tod54g, TT42, TT51, TT52a, TT52b, TT56a, TT57a, TT57b, TT06, TT12a, TT12b, Tod49a, Tod54b, Tod58c, TV99]. Matrix [New62, Tod53, Tod54c, NT58, TT91, Tod50b, Tod61c]. May [BW83, Cra87, Tod70a]. McGraw [Hou63, Tod56a]. McGraw-Hill [Hou63, Tod56a]. mean [Tod90b, TBDS92a, TBDS92b]. means [Tod78b, Tod88b]. memoriam [Bau07]. methods [Tod54e, Tod67a]. Metropolis [Ham65b, Ham65a, Ham65c, Jun64]. Mitteilungen [Tod52]. mixed [Tod78b, Tod88b]. modern [Mad12, TT47, Tod50c]. modification [Koj78]. Monte [Tod54e]. Motivation [Tod54f, Tod55c, Tod62c]. Mrs. [Tod54g].

N [Ham65b, Ham65a, Ham65c, Jun64]. N.F [Tod70c]. Natanson [Tod59a]. National [Tod59c, Tod87, Tod90a, Com36, Tod74, Tod75c]. Natur [Tod70c]. Nautical [Tod59b]. NBS [HT91]. NBS-INA [HT91]. Neumann [Tod74]. No [Tod70a, Tod52]. Nonsingular [Tau87]. North [BCEP94]. Notes [Tod50c]. Numbers [TT54, Tod51, Tod68, TT40a, TT56b]. Numeric [Cra87]. Numerical [Col64, HT91, Hou63, Olv63, Pra63, Ran80, Row79, Tod62d, Tod67a, Tod75c, Tod77, Tod78a, Too63, Alb07, Mad12, Tod50c, Tod54f, Tod55a, Tod55c, Tod56b, Tod57a, Tod57b, Tod59d, Tod62a, Tod62c, Fla64, Tod56a, Tod64]. numerischer [Tod55a, Tod57a, Tod57b].

O [Tod61a]. O. [Alz91]. Oberwolfach [BW83, Tod83]. Obituary [Tod54a]. Office [Tod59c, Tod59b]. Optimal [Tod67c, GT67, GT67]. Ordinary [Tod64]. orthogonal [Tod50d]. Ossicini [Tod71].

P [Tod59a, Tod61b, Hou63, Tod59c]. papers [Cra87, MTTT63c]. Parameter [Tod70c]. Parameters [Tod67c]. Partial [Tod64, Tod56b]. period [TT40c, TT41b]. periods [Tod90b]. Physical [Tod59c]. Pi [BBB00]. Piccard [Tod40]. Pioneer [WH97, Alb07]. points [Tod40]. polygons [FTT55a]. polynomial [Tod86, Tod61b]. Polynomialrecurrencology [MGT^{+}96]. Polynomials [KR79, CT83b, Tod59d]. powers [TT42]. pp [Fla64, Tod52, Tod56a, Tod70a]. Practical [ET46]. prehistory [Tod87, Tod90a]. presented [Cra87]. Press [Fla64, Tod70a]. Previous [TS85]. Price [Hou63]. Princeton [Cra87]. problem [CT83b, Tod49b, Tod56b, Tod65b]. Problems [BT48, BT50, DFT49, NST52, RKB^{+}87, RT87, RTDH90, Tod61c]. Proceedings [MTTT63a, MTTT63b, Tod56a, Tod70a, Cra87, BCEP94, Jun64]. processes [GT67, GT67]. products [TT51]. Professor [Ano97]. Programmgesteuerte [Tod52]. programs [NT58]. proof [FTT55a]. proofs [Lun94]. Properties [QZL^{+}10]. Pseudo [TT54, TT56b]. Pseudo-Random [TT54, TT56b]. Publication [Tod70a]. Publications [Olv63].

Quadrature [Tod71]. quaternions [TT40b]. Queries [Tod93].

R [Tod61b]. radius [KT58, KT59b, KT59a, KT60]. Raleigh [BCEP94]. Random [TT54, TT56b]. rate [GT67, GT67]. Rational [Tod51, TT51]. Reason [Tod55a, Tod57a, Tod57b]. Rechengeräte [Tod52]. Rechenmaschinen [Tod52]. Recurrence [Tod50e, HT73, Tod50c]. reducible [CT49]. relation [HT73]. Relations [Tod50e, Tod49b, Tod50c]. relative [Tod50d]. remarks [Tau87].

Reminiscences [Tod94b]. Reply [Tod54g]. Research [MTTT63c, Tod70a]. Response [TS85]. Review [Col64, Fla64, Ham65b, Ham65a, Ham65c, Hou63, Jun64, Mil65, Pra63, Ran80, Row79, Tod38, Tod40, Tod50a, Tod52, Tod56a, Tod58b, Tod58a, Tod59a, Tod59c, Tod59b, Tod61a, Tod61b, Tod63a, Tod64, Tod67a, Tod70a, Tod71, TBDS92a, TBDS92b, Too63]. Richardson [Tod54a]. Riemann [ST51]. root [HT73]. roots [CT36, TT51]. Ruthishauser [Tod52].

S [Tod38]. Saks [Tod38]. Schoenberg [Tod70a]. Scientific [Com36, Cra87, Nas90]. Sechsstellige [Tod70c]. Segments [Tod54c]. Selected [MTTT63c]. Self [TS85]. Self-Study [TS85]. sequences [Tod88b]. Service [ST46, ST46, TS47]. smallest [Tod65a]. Society [MTTT63a, MTTT63b]. Solution [Tau54, Tod50e, Tod64, Tod54e, TW55, Tod56b, Tod50c]. Solutions [BT48, BT50, DFT49, NST52, RKB^{+}87, RT87, RTDH90]. Some [ST46, TT47, TT60, Tod80a]. Sophie [Tod40]. source [BBB00]. Special [Tod59d, Tod70a]. Speed [Ham65b, Ham65a, Ham65c, Jun64, MTTT63c, MTTT63a, MTTT63b]. Speiser [Tod52]. Spline [Tod70a]. square [HT73]. stability [Tod56b]. Stable [KR79]. Standards [Tod75c, Tod87, Tod90a]. Stationery [Tod59c]. Statistics [Tod70b]. Stiefel [Tod52]. Stieltjes [LT72, LT73]. Study [TS85]. Subtabulation [Tod59b]. sum [ST81]. sums [ST51]. Superpositions [Tod35, Tod36]. Survey [Col64, Hou63, Olv63, Pra63, Tod62d, Too63]. Swiss [Tod52]. symmetric [TT06]. Symposia [MTTT63c, Tod56a]. Symposium [Jun64, MTTT63a, MTTT63b, Tod70a]. Systems [TTT57, TT52a, TT52b, Tau54, TT57b, TT12b, Tod54g, TT57a, TT12a].

Table [Tod51, Tod70c]. Tables [Tod59c, Tod59c, Tod59b, Tod58a]. Tafel [Tod70c]. tangent [Tod49b]. Tarski [Tod50a]. Taub [Ham65b, Ham65a, Ham65c, Jun64]. Taussky [Alz91, Lun94, Yin94]. Testing [TT54, TT56b]. theorem [Tod37]. theorems [TT48a, TT48b, TT49]. Theory [Fla64, Mil65, Tod59a, Tod63b, Tod62b, Tod84, Tod38]. Time [KR79, ST46]. Todd [Coh97, Fla64, Ham65b, Ham65a, Ham65c, Hou63, Jun64, Row79, Alb07, Alz91, Ano97, Bau07, Col64, Koj78, Lun94, Mad12, Mil65, OR07, Olv63, Pra63, Ran80, Tau87, Too63, WH97, Yin94]. Toeplitz [TT06]. Tompkins [Jun64, Ham65b, Ham65a, Ham65c]. Transfinite [Tod35, Tod36]. transformation [Tod84]. transformations [TT56a]. Tree [JB57, JB12]. triangular [TT06]. triangulation [Koj78]. trigonometrical [ST81]. two [Alz91, KT76]. Type [Tod58a].

U.S [Tod70a]. U.S. [Tod87, Tod90a]. UCLA [HT91]. ultrametric [TV99]. univalence [KT58, KT59b, KT59a, KT60]. University [Tod70a]. use [Tod59c].

v [Tod59c]. Values [Tod58a]. variable [TT60]. VI [Tod56a]. Vol [Row79, Tod56a]. Volume [Ran80, Tod77, Tod78a]. Vuorinen [TBDS92a, TBDS92b].

W. [Tod67b]. Walther [Tod93]. War [ST46]. War-Time [ST46]. Weierstrass [Tod90b]. Weighted [QZL^{+}10]. Whereabouts [Ano50]. Wirtinger [FTT55b]. Wis [Tod70a]. Wisconsin [Tod70a]. Wiss [Tod70c]. working [Tod54f, Tod55c, Tod62c].

xii [Tod70a]. xvi [Hou63].

York [Fla64, Tod56a, Tod70a].

Zolotareff [Tod67b]. Zolotarev [CT83b, Tod84, Tod88a].

References

Albers:2007:JTN

[Alb07]  Don Albers.  John Todd—

numerical mathematics pioneer. College Mathematics Journal, 38 (1):2–23, 2007. ISSN 0746-8342,1931-1346.

Alzer:1991:CTI

[Alz91] Horst Alzer. Converses of two inequalities of Ky Fan, O. Taussky, and J. Todd. Journal of Mathematical Analysis and Applications, 161(1):142–147, 1991. CO-DEN JMANAK. ISSN 0022-247X,1096-0813.

Anonymous:1950:W

[Ano50]  Anonymous.  Whereabouts.

Physics Today, 3(3):37, March 1950. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). From the article: “Two British mathematicians, Mr. and Mrs. John Todd, have joined the staff of the National Bureau of Standards’ applied mathematics laboratories. Mr. Todd, formerly of Kings College, London University, and Queens University in Belfast, has been appointed chief of the Bureau’s computation laboratory, where he will be in charge of mathematical research as well as of the computing services offered by the laboratory to various government agencies. Mrs. Todd (Dr.

Olga Taussky), as a technical consultant, will do fundamental research in problems of mathematical analysis, algebra, and number theory in their relation to the use of high-speed computing machines”.

Anonymous:1997:PJT

[Ano97] Anonymous. Professor John Todd. The Telegraph, ??(??): ??, July 5, 1997. URL https:

//www.telegraph.co.uk/news/
obituaries/1556507/Professor-
John-Todd.html.

Bauer:2007:MJT

[Bau07] Friedrich L. Bauer. In memoriam John Todd (1911–2007). Numerische Mathematik, 108(1):1– 6, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berggren:2000:PSB

[BBB00] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-98946-3 (hardcover). xx + 736 pp. LCCN QA484 .P5 2000.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Boas:1948:PSE

[BT48] R. P. Boas, Jr. and Olga Taussky (Mrs. John Todd). Problems and solutions: Elementary problems: Solutions: E775. American Mathematical Monthly, 55 (2):99, February 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/
stable/2305751. See also [?].

Browne:1950:APS

[BT50] D. H. Browne and John Todd. Advanced problems and solutions: Solutions: 4341. American Mathematical Monthly, 57(8): 568–569, October 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/
stable/2307960. See also [?].

Beckenbach:1983:GII

[BW83] Edwin F. Beckenbach and Wolfgang Walter, editors. General Inequalities 3: 3rd International Conference on General Inequalities Oberwolfach, April 26– May 2, 1981, volume 64 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1983. ISBN 3-7643-1539-3. LCCN QA295 .I57 1981. URL https:/

/link.springer.com/book/10.
1007/978-3-0348-6290-5.

Cohen:1997:IJT

[Coh97] Shirley K. Cohen. Interview with John Todd, March 29 and April 5, 1996. California Institute of Technology Archives, 1997.

URL https://www.jstor.org/
stable/community.30990609.

Cole:1964:BRS

[Col64] A. J. Cole. Book review: A Survey of Numerical Analysis by John Todd. Mathematical Gazette, 48(364):247, May 1964. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.
jstor.org/stable/3613594.

Comrie:1936:IIS

[Com36] L. J. Comrie. Inverse interpolation and scientific applications of the National Accounting Machine. Supplement to the Journal of the Royal Statistical Society, 3(2):87–114, ???? 1936. CODEN ???? ISSN 1466-6162.

URL http://www.jstor.org/
stable/2983666. See [Tod74].

Crane:1987:HAC

[Cra87] G. E. Crane, editor. HSNC’87:

ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Cooper:1936:LR

[CT36] Ralph Cooper and John Todd. The large roots of cos z = az + c. Philosophical Magazine Series 7 (1926–1955), 21:249–262, 1936. CODEN PHMAA4. ISSN 0031-8086.

Chowla:1949:DRI

[CT49] S. D. Chowla and John Todd. The density of reducible integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1:297–299, 1949. CODEN CJMAAB. ISSN 0008-414X,1496-4279.

Carlson:1983:DBE

[CT83a] B. C. Carlson and John Todd. The degenerating behavior of elliptic functions. SIAM Journal on Numerical Analysis, 20(6): 1120–1129, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2157146.

Carlson:1983:ZFP

[CT83b] B. C. Carlson and John Todd. Zolotarev’s first problem — the best approximation by polynomials of degree ≤ n − 2 to x^{n} − nσx^{n−1} in [−1, 1]. Aequationes Mathematicae, 26:1– 33, 1983. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic).

Dubisch:1949:EPSb

[DFT49] Roy Dubisch, Jacob Feldman, and John Todd. Elementary problems and solutions: E852. American Mathematical Monthly, 56(8):554– 555, October 1949. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/
stable/2305537. See also [?].

Erdelyi:1946:AIP

[ET46] A. Erdélyi and John Todd. Advanced instruction in practical mathematics. Nature, 158(4020): 690–692, November 16, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Flanders:1964:BRN

[Fla64] Harley Flanders. Book review:

Numerical Mathematics: Introduction to the Constructive Theory of Functions, John Todd. Academic Press, New York, 1963. 127 pp. Illus. $6. Science, 144(3617):402, April 24, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL https://www.
jstor.org/stable/1713434.

Fan:1955:DI

[FT55] Ky Fan and John Todd. A determinantal inequality. Journal of the London Mathematical Society, 30(1):58–64, January 1955. CODEN JLMSAK. ISSN 0024-6107,1469-7750.

Fan:1955:API

[FTT55a] Ky Fan, Olga Taussky, and John Todd. An algebraic proof of the isoperimetric inequality for polygons. Journal of the Washington Academy of Sciences, 45:339–342, 1955. CO-DEN JWASA3. ISSN 0043-0439.

                                        URL https://www.jstor.org/
                                        stable/24534441.

Fan:1955:DAI

[FTT55b] Ky Fan, Olga Taussky, and John Todd. Discrete analogs of inequalities of Wirtinger. Monatshefte für Mathematik, 59:73–90, 1955. CODEN MNMTA2. ISSN 0026-9255,1436-5081.

Gaier:1967:RCO

[GT67] Dieter Gaier and John Todd. On the rate of convergence of optimal ADI processes. Numerische Mathematik, 9(5):452– 459, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gaier:1966:RCO

[GT67] Dieter Gaier and John Todd. On the rate of convergence of optimal ADI processes. Numerische Mathematik, 9:452–459, 1966/67. CODEN NUMMA7. ISSN 0029-599X,0945-3245.

Hammer:1965:BRBb

[Ham65a] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:/

                                        /links.jstor.org/sici?sici=

0040-1706%28196502%297%3A1%
3C82%3AEAHSCA%3E2.0.CO%3B2-
9.

Hammer:1965:BRBa

[Ham65b] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis; A. H. Taub; John Todd; C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.
jstor.org/stable/1266139.

Hammer:1965:BRE

[Ham65c] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:/

/links.jstor.org/sici?sici=
0040-1706%28196502%297%3A1%
3C82%3AEAHSCA%3E2.0.CO%3B2-
9.

Householder:1963:BRJ

[Hou63] A. S. Householder. Book review: John Todd, Editor, A Survey of Numerical Analysis, McGraw-Hill Book Co., 1962, 23.5 cm., xvi + 589 p. Price $12.50. Mathematics of Computation, 17(81):89, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.

                                        org/journals/mcom/1963-17-
                                        081/S0025-5718-1963-1781088-
                                        8;   http://www.ams.org/journals/
                                        mcom/1963-17-081/S0025-5718-
                                        1963-1781088-8/S0025-5718-
                                        1963-1781088-8.pdf;   https:
                                        //www.jstor.org/stable/2003741.

Hwang:1973:RRS

[HT73] W. G. Hwang and John Todd. A recurrence relation for the square root. Journal of Approximation Theory, 9:299–306, 1973. CO-DEN JAXTAZ. ISSN 0021-9045,1096-0430.

Hestenes:1991:NIN

[HT91] Magnus Rudolph Hestenes and John Todd. NBS-INA, the Institute for Numerical Analysis, UCLA 1947–1954. National Technical Information Service, Washington, DC, USA, August 1991. various pp. Cover title: Mathematicians learning to use computers: the Institute for Numerical Analysis, UCLA, 1947– 1954. Spine title: The Institute for Numerical Analysis, UCLA, 1947–1954. Sponsored in part by the Mathematical Association of America. Shipping list no.: 92-1306-M.

James:1957:TM

[JB57] Glenn James and E. F. Beckenbach, editors. The Tree of Mathematics. Digest Press, Pacoima, CA, USA, 1957. xvii + 403 pp. LCCN QA37 .J26.

James:2012:TM

[JB12] Glenn James and E. F. Beckenbach, editors. The Tree of Mathematics. Digest Press, Literary Licensing, LLC, 2012. ISBN 1-258-28562-2. xvii + 403 pp. LCCN QA37 .J26.

Juncosa:1964:BRB

[Jun64] Mario Juncosa. Book review:

Proceedings of the 15th Symposium in Applied Mathematics of the A.M.S. Experimental Arithmetic, High Speed Computing and Mathematics (N. C. Metropolis, A. H. Taub, John Todd C. B. Tompkins). SIAM Review, 6(4):468–471, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL https://www.
jstor.org/stable/2027983.

Kojima:1978:MTT

[Koj78] Masakazu Kojima. A modification of Todd’s triangulation J_{3}. Mathematical Programming, 15 (2):223–227, 1978. CODEN MH-PGA4. ISSN 0025-5610,1436-4646.

Kusterer:1979:SEP

[KR79] Roland Kusterer and Manfred Reimer. Stable evaluation of polynomials in time log n. Mathematics of Computation, 33 (147):1019–1031, July 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kreyszig:1958:RUE

[KT58] Erwin Kreyszig and John Todd.

The radius of univalence of the error function. Bulletin of the American Mathematical Society, 64:363–364, 1958. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Kreyszig:1959:RUF

[KT59a] Erwin Kreyszig and John Todd. On the radius of univalence of the function exp z^{2} ∫_{0}^{z} exp(−t^{2}) dt. Pacific Journal of Mathematics, 9:123–127, 1959. CO-DEN PJMAAI. ISSN 0030-8730,1945-5844. URL http:

                                        //projecteuclid.org/euclid.
                                        pjm/1103039457.

Kreyszig:1959:RUE

[KT59b] Erwin Kreyszig and John Todd. The radius of univalence of the error function. Numerische Mathematik, 1:78–89, 1959. CO-DEN NUMMA7. ISSN 0029-599X,0945-3245.

Kreyszig:1960:RUB

[KT60] Erwin Kreyszig and John Todd. The radius of univalence of Bessel functions. I. Illinois Journal of Mathematics, 4:143– 149, 1960. CODEN IJMTAW. ISSN 0019-2082 (print), 1945-6581 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        ijm/1255455740.

Kreyszig:1976:TD

[KT76] E. Kreyszig and J. Todd. QR in two dimensions. Elemente der Mathematik, 31:109–114, 1976. CODEN ELMMAF. ISSN 0013-6018 (print), 1420-8962 (electronic).

Liang:1972:SC

[LT72] J. J. Y. Liang and John Todd.

The Stieltjes constants. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 76B:161–178, 1972. CO-DEN JNBBAU. ISSN 0022-4340.

Liang:1973:SC

[LT73] J. J. Y. Liang and John Todd.

The Stieltjes constants. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 76:161–178, 1973. CO-DEN JNBBAU. ISSN 0022-4340 (print), 2376-5283 (electronic).

Lunter:1994:NPG

[Lun94] Gerton Lunter. New proofs and a generalization of inequalities of Fan, Taussky, and Todd. Journal of Mathematical Analysis and Applications, 185(2): 464–476, July 15, 1994. CO-DEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Madden:2012:JTD

[Mad12] Niall Madden. John Todd and the development of modern numerical analysis. Irish Mathematical Society Bulletin, 69: 11–24, Summer 2012. ISSN 0791-5578. URL https:/

/www.maths.tcd.ie/pub/ims/
bull69/Madden.pdf.

Mason:1996:P

[MGT^{+}96] Eric H. Mason, Stephen M.

Gagola, Jr., John Todd, Theodore J.

Rivlin, and David Callan. Polynomialrecurrencology: 10300. American Mathematical Monthly, 103(3):272–273, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.
jstor.org/stable/2975388.

Miller:1965:BRI

[Mil65] G. F. Miller. Book review: Introduction to the Constructive Theory of Functions by John Todd. Mathematical Gazette, 49 (370):486, December 1965. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL https://www.jstor.org/
stable/3612241.

Metropolis:1963:EAH

[MTTT63a] N. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins, editors. Experimental arithmetic, high speed computing and mathematics: Proceedings of the fifteenth Symposium in Applied Mathematics of the American Mathematical Society held in Chicago, Illinois, April 12– 14, 1962 and Atlantic City, New Jersey, April 16-19, 1962, volume 15. American Mathematical Society, Providence, RI, USA, 1963. LCCN QA297 .S987 1962.

Metropolis:1963:PFS

[MTTT63b] N. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins, editors. Experimental arithmetic, high speed computing and mathematics: Proceedings of the fifteenth Symposium in Applied Mathematics of the American Mathematical Society held in Chicago, Illinois, April 12– 14, 1962 and Atlantic City, New Jersey, April 16-19, 1962, volume 15. American Mathematical Society, Providence, RI, USA, 1963. ISSN 0160-7634. LCCN QA297 .S987 1962.

Metropolis:1963:IBM

[MTTT63c] N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins, editors. Interactions between Mathematical Research and High-Speed Computing: Symposia: Selected Papers. American Mathematical Society, Providence, RI, USA, 1963. LCCN ????

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Newman:1962:MC

[New62] Morris Newman. Matrix computations. In John Todd, editor, Survey of Numerical Analysis, pages 222–254. McGraw-Hill, New York, NY, USA, 1962.

Newman:1952:APSd

[NST52] D. J. Newman, J. D. Swift, and John Todd. Advanced problems and solutions: 4436. American Mathematical Monthly, 59(9):

640–641, November 1952. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/2306782. See

also [?, ?].

Newman:1958:EMI

[NT58] Morris Newman and John Todd. The evaluation of matrix inversion programs. Journal of the Society for Industrial and Applied Mathematics, 6(4):466– 476, December 1958. CO-DEN JSIMAV. ISSN 0368-4245.

                                        URL https://www.jstor.org/
                                        stable/2098717.

Newman:1962:AC

[NT62] Morris Newman and John Todd. Automatic computers. In Survey of numerical analysis, pages 160– 207. McGraw-Hill, New York, NY, USA, 1962.

Olver:1963:RPS

[Olv63] F. W. J. Olver. Recent publications: A Survey of Numerical Analysis, by John Todd. American Mathematical Monthly, 70 (2):224–225, February 1963. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/2312919.

OConnor:2007:JT

[OR07] J. J. O’Connor and E. F. Robertson. Jack Todd. St. Andrews MacTutor Web site, August 2007. URL https://

                                        mathshistory.st-andrews.ac.
                                        uk/Biographies/Todd_John/.

Ostrowski:1949:B

[OTT49] Alexander M. Ostrowski, Olga Taussky Todd, and John Todd. [Bibliography]. Report ????, U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1949. ???? pp.

Prager:1963:BRS

[Pra63] W. Prager. Book review: A Survey of Numerical Analysis by John Todd. Quarterly of Applied Mathematics, 20(4):396, January 1963. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic). URL https://www.

jstor.org/stable/43636451.

Qiao:2010:WDW

[QZL^{+}10] Xingye Qiao, Hao Helen Zhang, Yufeng Liu, Michael J. Todd, and J. S. Marron. Weighted distance weighted discrimination and its asymptotic properties. Journal of the American Statistical Association, 105(489):401– 414, March 2010. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Randall:1980:BRB

[Ran80] T. J. Randall. Book review:

Basic Numerical Mathematics; Volume 1, Numerical Analysis, and Volume 2, Numerical Algebra by John Todd. Mathematical Gazette, 64(430):297– 298, December 1980. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL https://www.jstor.org/
stable/3616749.

Rawsthorne:1987:PSE

[RKB^{+}87] Daniel Rawsthorne, Erwin Kronheimer, P. A. Batnik, Alexandru Lupaş, William F. Eberlein, Bruce Reznick, and John Todd. Problems and solutions: Elementary problems: E3219– E3224. American Mathematical Monthly, 94(7):680–681, August/September 1987. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/
stable/2322225. See also [?, ?,

?, ?, ?, ?, RTDH90].

Rowland:1979:BRB

[Row79] John H. Rowland. Book review: Basic Numerical Mathematics, Vol. 2: Numerical Algebra (John Todd). SIAM Review, 21(3):404, July 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://www.jstor.org/
stable/2029587.

Reznick:1987:PSE

[RT87] Bruce Reznick and John Todd. Problems and solutions: Elementary problems: E3224. American Mathematical Monthly, 94(7): 681, August/September 1987. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.
jstor.org/stable/2323923.

Reznick:1990:PSS

[RTDH90] Bruce Reznick, John Todd, Jim Delany, and Kempton Huehn. Problems and solutions: Solutions of elementary problems:

E3224. American Mathematical Monthly, 97(2):153–154, February 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RKB^{+}87].

Sadler:1946:MGS

[ST46] D. H. Sadler and John Todd.

Mathematics in government service and industry: Some deductions from the war-time experience of the Admiralty Computing Service. Nature, 157(3992): 571–573, May 4, 1946. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Szasz:1951:CCR

[ST51] Otto Szász and John Todd. Convergence of Cauchy–Riemann sums to Cauchy–Riemann integrals. Journal of Research of the National Bureau of Standards (1934), 47:191–196, 1951. ISSN 0091-0635.

Stembridge:1981:TS

[ST81] John R. Stembridge and John Todd. On a trigonometrical sum. Linear Algebra and its Applications, 35(??):287–291, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379581902809.

Taussky:1954:CSS

[Tau54] Olga Taussky, editor. Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, volume 39 of Applied Mathematics Series. U.S. National Bureau of Standards, Gaithersburg, MD, USA, September 30, 1954.

Taussky:1987:NCC

[Tau87]  Olga Taussky.  Nonsingular

cubic curves as determinantal loci. Historical remarks by John Todd. Journal of Mathematical and Physical Sciences, 21(6): 665–678, 1987. ISSN 0047-2557. Ramanujan memorial volume.

Todd:1992:BRB

[TBDS92a] John Todd, Bill Braden, Bernard Danloy, and Frank Schmidt. Book review: Inequality of the AGM and the Logarithmic Mean (B. C. Carlson and M. Vuorinen). SIAM Review, 34(4): 653–654, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Todd:1992:BRI

[TBDS92b] John Todd, Bill Braden, Bernard Danloy, and Frank Schmidt. Book review: Inequality of the AGM and the Logarithmic Mean (B. C. Carlson and M. Vuorinen). SIAM Review, 34(4): 653–654, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Todd:1935:SFT

[Tod35] John Todd. Superpositions of functions. I: Transfinite superpositions of absolutely continuous functions. Journal of the London Mathematical Society, 10(3):166– 171, July 1935. CODEN JLM-SAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Todd:1936:SFI

[Tod36] John Todd. Superpositions of functions. II. Transfinite superpositions of absolutely continuous functions. Proceedings of the London Mathematical Society. Second Series, 41(1):433– 439, 1936. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244x (electronic).

Todd:1937:TA

[Tod37] J. Todd. A theorem in arithmetic. Mathematical Gazette, 21:423–424, 1937. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL https://www.jstor.org/
stable/3605843.

Todd:1938:BRT

[Tod38] John Todd. Book review: Theory of the Integral by S. Saks. Mathematical Gazette, 22(248): 84, February 1938. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL https://www.jstor.org/
stable/3607456.

Todd:1940:BRE

[Tod40] John Todd. Book review: Sur les ensembles de distances des ensembles de points d’un espace euclidien by Sophie Piccard. Mathematical Gazette, 24 (261):302, October 1940. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL https://www.jstor.org/
stable/3605468.

Todd:1949:CCM

[Tod49a] John Todd. The condition of certain matrices. I. Quarterly Journal of Mechanics and Applied Mathematics, 2:469–472, 1949. CODEN QJMMAV. ISSN 0033-5614,1464-3855.

Todd:1949:PAT

[Tod49b] John Todd. A problem on arc tangent relations. American Mathematical Monthly, 56: 517–528, 1949. CODEN AM-MYAE. ISSN 0002-9890,1930-

                                        0972.  URL https://www.
                                        jstor.org/stable/2305526.

Todd:1950:BRC

[Tod50a] John Todd. Book review: Cardinal Algebras by Alfred Tarski, Bjarni Jónsson. Science Progress, 38(149):152, Jan-

                                        uary 1950. URL https://www.
                                        jstor.org/stable/43413694.

Todd:1950:CCM

[Tod50b] John Todd. The condition of a certain matrix. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 46:116–118, 1950. CO-DEN PCPSA4. ISSN 0008-1981.

Todd:1950:NMN

[Tod50c] John Todd. Notes on modern numerical analysis. I. Solution of differential equations by recurrence relations. Mathematical Tables and Other Aids to Computation, 4:39–44, 1950. CO-DEN MTTCAS. ISSN 0891-6837,2326-4853.

Todd:1950:REL

[Tod50d] John Todd. On the relative extrema of the Laguerre orthogonal functions. Bollettino della Unione Matematica Italiana. Series III, 5:122–125, 1950.

Todd:1950:SDE

[Tod50e] John Todd. Solution of differential equations by recurrence relations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 4(29):39– 44, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

URL https://www.jstor.org/
stable/2002701.

Todd:1951:TAR

[Tod51] John Todd. Table of Arctangents of Rational Numbers, volume 11 of National Bureau of Standards Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1951. xi + 105 pp.

Todd:1952:BRP

[Tod52] John Todd. Book review: Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen), By H. Ruthishauser, A. Speiser, and E. Stiefel. (Mitteilungen aus dem Institut fur angewandte Mathematik, no. 2.) Basel, Birkhäuser, 1951. 102 pp. 8.50 Swiss fr. Bulletin of the American Mathematical Society, 58(2):278–279, March 1952. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Todd:1953:EIM

[Tod53] John Todd. Experiments on the inversion of a 16 × 16 matrix. In L. J. (Lowell J.) Paige and Olga Taussky, editors, Simultaneous Linear Equations and the Determination of Eigenvalues, number 29 in National Bureau of Standards Applied Mathematics Series, pages 113–115. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1953.

Todd:1954:OFR

[Tod54a] J. Todd. Obituary: L. F. Richardson (1881–1953). Mathematical Tables and Other Aids to Computation, 8:242–245, 1954. CODEN MTTCAS. ISSN 0891-6837,2326-4853.

Todd:1954:CCM

[Tod54b] John Todd. The condition of certain matrices. II. Archiv der Mathematik, 5:249–257, 1954. CODEN ACVMAL. ISSN 0003-889X,1420-8938.

Todd:1954:CFS

[Tod54c] John Todd. The condition of the finite segments of the Hilbert matrix. In Olga Taussky, editor, Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, number 39 in National Bureau of Standards Applied Mathematics Series, pages 109–116. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1954.

Todd:1954:EEI

[Tod54d] John Todd. Evaluation of the exponential integral for large complex arguments. Journal of Research of the National Bureau of Standards (1934), 52:313–317, 1954. ISSN 0091-0635.

Todd:1954:ESD

[Tod54e] John Todd. Experiments in the solution of differential equations by Monte Carlo methods. Journal of the Washington Academy of Sciences, 44:377–381, 1954. CODEN JWASA3. ISSN 0043-

0439.  URL https://www.
jstor.org/stable/24533326.

Todd:1954:MWN

[Tod54f] John Todd. Motivation for working on numerical analysis. In ????, editor, Transactions of 2nd Symposium on Applied Mathematics, 29 April 1954, University of Chicago, page ???? American Mathematical Society, Providence, RI, USA, 1954.

Todd:1954:RMB

[Tod54g] John Todd. Reply to Mrs. Bernice Brown’s Comment on “Systems of Equations, Matrices, and Determinants”. Mathematics Magazine, 27(5):288, May/June 1954. CODEN MAMGA8. ISSN 0025-570X. URL https://www.

jstor.org/stable/3029242.

Todd:1955:BBN

[Tod55a] John Todd. Begründung für die Beschäftigung mit numerischer Analysis. (German) [Reason for engaging in numerical analysis]. Jber. Deutsch. Math.–Verein, 58: 11–38, 1955. ISSN 0012-0456 (print), 1869-7135 (electronic).

Todd:1955:ECC

[Tod55b] John Todd. Experiments in the computation of conformal maps. Technical Report 42, United States Government Printing Office, Washington, DC, USA, 1955. 61 pp.

Todd:1955:MWN

[Tod55c] John Todd. Motivation for working in numerical analysis. Communications on Pure and Applied Mathematics (New York), 8(1): 97–116, February 1955. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Todd:1956:BRN

[Tod56a] John Todd. Book review: Numerical Analysis. Proceedings of Symposia in Applied Mathematics, vol. VI. John H. Curtiss, Ed. McGraw-Hill, New York, 1956. 303 pp. $9.75. Science, 124 (3233):1211, December 14, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL https://www.
jstor.org/stable/1752848.

Todd:1956:DAP

[Tod56b] John Todd. A direct approach to the problem of stability in the numerical solution of partial differential equations. Communications on Pure and Applied Mathematics (New York), 9(3): 597–612, 1956. CODEN CPA-MAT, CPMAMV. ISSN 0010-3640,1097-0312.

Todd:1957:BBNa

[Tod57a] J. Todd. Begründung für die Beschäftigung mit numerischer Analysis. (German) [Reason for engaging in numerical analysis]. Mat. Prosveshchenie, 1(??):75– 86, 1957.

Todd:1957:BBNb

[Tod57b] J. Todd. Begründung für die Beschäftigung mit numerischer Analysis. (German) [Reason for engaging in numerical analysis]. Mat. Prosveshchenie, 2(??):97– 110, 1957.

Todd:1958:BRT

[Tod58a] John Todd. Book review:

Tables of Values of 16 Integrals of Algebraic-Hyperbolic Type by Chih-Bing Ling. Mathematics of Computation, 12(64): 306, October 1958. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.jstor.org/
stable/2002419.

Todd:1958:BRC

[Tod58b] John Todd. Book review: Cornelius Lanczos, Applied analysis. Bulletin of the American Mathematical Society, 64(4):210–211, ???? 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http:

//projecteuclid.org/euclid.
bams/1183522571.

Todd:1958:CCM

[Tod58c] John Todd. The condition of certain matrices. III. Journal of Research of the National Bureau of Standards (1934), 60:1–7, 1958. ISSN 0160-1741.

Todd:1959:BRC

[Tod59a] John Todd. Book review: Constructive Theory of Function by I. P. Natanson, K. Bogel. Mathematics of Computation, 13(65): 64–67, January 1959. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.jstor.org/
stable/2002205.

Todd:1959:BRS

[Tod59b] John Todd. Book review: Subtabulation, A Companion Booklet to Interpolation and Allied Tables, by Her Majesty’s Nautical Almanac Office. Mathematics of Computation, 13(66): 127–129, April 1959. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.jstor.org/
stable/2001970.

Todd:1959:BRF

[Tod59c] John Todd. Book review: L. Fox. The use and construction of mathematical tables. National Physical Lab., Math. Tables, v. 1, Her Majesty’s Stationery Office, London, England, 1956, 60 p., 27.5 cm.;. Mathematics of Computation, 13(65):61– 64, January 1959. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.jstor.org/
stable/2002204.

Todd:1959:SPN

[Tod59d] John Todd. Special polynomials in numerical analysis. In On numerical approximation. Proceedings of a Symposium, Madison, April 21-23, 1958, volume no. 1 of Publication of the Mathematics Research Center, U.S. Army, the University of Wisconsin, pages 423–446. University of Wisconsin Press, Madison, WI, USA, 1959. Edited by R. E. Langer.

Todd:1960:CCI

[Tod60]  John Todd.  The condition

of certain integral equations. In Symposium on the numerical treatment of ordinary differential equations, integral and integro-differential equations (Rome, 1960), pages 306–311. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1960.

Todd:1961:BRD

[Tod61a] John Todd. Book review: Differenzenrechnung by A. O. Gelfond. Mathematics of Computation, 15(74):202–203, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.
jstor.org/stable/2004236.

Todd:1961:BRP

[Tod61b] John Todd. Book review:

Polynomial Expansions of Analytic Functions by R. P. Boas, Jr., R. C. Buck. Mathematics of Computation, 15(73):104– 105, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2003118.

Todd:1961:CPC

[Tod61c] John Todd. Computational problems concerning the Hilbert matrix. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 65B: 19–22, 1961. CODEN JNBBAU. ISSN 0022-4340.

Todd:1962:CNA

[Tod62a] John Todd. Classical numerical analysis. In Survey of numerical analysis, pages 27–118. McGraw-Hill, New York, NY, USA, 1962.

Todd:1962:CTF

[Tod62b] John Todd. The constructive theory of functions. In Survey of numerical analysis, pages 119– 159. McGraw-Hill, New York, NY, USA, 1962.

Todd:1962:MWN

[Tod62c] John Todd. Motivation for working in numerical analysis. In Survey of numerical analysis, pages 1–26. McGraw-Hill, New York, NY, USA, 1962.

Todd:1962:SNA

[Tod62d] John Todd, editor. Survey of Numerical Analysis. Mc-Graw-Hill, New York, NY, USA, 1962. xvi + 589 pp. LCCN QA297 .T63. URL https:/

/babel.hathitrust.org/cgi/
pt?id=mdp.39015000983901.

Todd:1963:BRA

[Tod63a] John Todd. Book review: Advances in Computers by Franz L. Alt. Mathematics of Computation, 17(81):98–99, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

jstor.org/stable/2003753.

Todd:1963:ICT

[Tod63b] John Todd. Introduction to the Constructive Theory of Functions. ISNM. International Series of Numerical Mathematics. Academic Press, New York, USA, 1963. ISSN 0373-3149. 127 pp.

Todd:1964:BRN

[Tod64] John Todd. Book review: Numerical Solution of Ordinary and Partial Differential Equations by L. Fox. American Scientist, 52(1):116A–117A, March 1964. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL https://www.
jstor.org/stable/27838969.

Todd:1965:SIG

[Tod65a] John Todd. On smallest isolated Gerschgorin disks for eigenvalues. Numerische Mathematik, 7:171–175, 1965. CO-DEN NUMMA7. ISSN 0029-599X,0945-3245.

Todd:1965:PED

[Tod65b] John Todd. The problem of error in digital computation. In Error in Digital Computation, Vol. 1 (Proceedings of an Advanced Seminar Conducted by [the] Mathematics Research Center, U.S. Army, University of Wisconsin, Madison, Wisconsin, 1964), pages 3–41. Wiley, New York, NY, USA, 1965.

Todd:1967:BRA

[Tod67a] John Todd. Book review: Analysis of Numerical Methods by E. Isaacson, H. B. Keller. American Scientist, 55(2):240A– 240A, June 1967. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL https://www.jstor.org/
stable/27836902.

Todd:1967:ICZ

[Tod67b] John Todd. Inequalities of Chebyshev, Zolotareff, Cauer, and W. B. Jordan. In Inequalities (Proceedings of a Symposium at Wright-Patterson Air Force Base, Ohio, 1965), pages 321– 328. Academic Press, New York, USA, 1967.

Todd:1967:OP

[Tod67c] John Todd. Optimal ADIparameters. In Funktionalanalysis, Approximationstheorie, Numerische Mathematik (Oberwolfach, 1965), volume 7 of Internat. Schriftenreihe Numer. Math., pages 58–69. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1967.

Todd:1968:CN

[Tod68] J. Todd. On condition numbers. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7 (number 165) of Éditions Centre Nat. Recherche Sci., Paris, pages 141– 159. CNRS, Paris, France, 1968.

Todd:1970:BRA

[Tod70a] J. Todd. Book review: Approximations with Special Emphasis on Spline Functions. Proceedings of a symposium, Madison, Wis., May 1969. I. J. Schoenberg, Ed. Academic Press, New York, 1969. xii + 492 pp., illus. $10. Publication No. 23 of the U.S. Army Mathematics Research Center, University of Wisconsin. Science, 168 (3936):1196, June 5, 1970. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL https://www.jstor.org/
stable/1729620.

Todd:1970:CS

[Tod70b] John Todd. Comprehensive statistics. Critical Survey, 5(1):49–53, Winter 1970.

URL https://www.jstor.org/
stable/41553850.

Todd:1970:TES

[Tod70c] John Todd. Table erratum:

“Sechsstellige Tafel der Cauer-Parameter” (Bayer. Akad. Wiss. Math.-Natur. Kl. Abh. (N.F.) Heft 67 (1955)) by Ernst Glowatzki. Mathematics of Computation, 24(112):999, 1970. CO-DEN MCMPAF. ISSN 0025-5718,1088-6842. URL http:/

                                        /links.jstor.org/sici?sici=
                                        0025-5718(197010)24:112<999:
                                        TE>2.0.CO;2-X&origin=MSN.

Todd:1971:BRQ

[Tod71]  John Todd.  Book review:

Quadrature Formulae by A. Ghizzetti, A. Ossicini. American Scientist, 59(3):380, May/June 1971. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL https://www.

                                        jstor.org/stable/27829677.

Todd:1974:JNN

[Tod74] John Todd. John von Neumann and the National Accounting Machine. SIAM Review, 16 (4):526–530, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2028693.  See also

[Com36].

Todd:1975:CLC

[Tod75a] John Todd. Corrigendum: “The Lemniscate Constants”. Communications of the ACM, 18 (8):462, August 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Tod75b].

Todd:1975:LC

[Tod75b] John Todd. The lemniscate constants. Communications of the ACM, 18(1):14–19, January 1975. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See corrigendum [Tod75a]. Reprinted in [Tod00].

Todd:1975:NAN

[Tod75c] John Todd. Numerical analysis at the National Bureau of Standards. SIAM Review, 17 (2):361–370, April 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://www.jstor.org/
stable/2029181.

Todd:1977:BNMb

[Tod77] John Todd. Basic Numerical Mathematics, Volume 2: Numerical Algebra. Academic Press, New York, USA, 1977. ISBN 0-12-692402-3. ISSN 0373-3149. 216 pp. LCCN QA297 .T58 1978.

Todd:1978:BNMa

[Tod78a] John Todd. Basic Numerical Mathematics, Volume 1: Numerical Analysis, volume 14, 22 of International series of numerical mathematics. Academic Press, New York, USA, 1978. ISBN 0-12-692401-5. ISSN 0373-3149. 253 pp. LCCN QA297 .T58 1978.

Todd:1978:MLM

[Tod78b] John Todd. The many limits of mixed means, I. In International Series of Numerical Mathematics / Internationale Schriftenreihe zur Numerischen Mathematik / Série Internationale d’Analyse Numérique, volume 41, pages 5– 22. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1978. General Inequalities 1 / Allgemeine Ungleichungen. 1, Tag. Oberwolfach 1976.

Todd:1980:SAE

[Tod80a] J. Todd. Some applications of elliptic functions and integrals. In Numerical methods, Keszthely 1977, Colloq. Math. Soc. Janos Bolyai, volume 22, pages 591– 618. ????, ????, 1980.

Todd:1980:JHC

[Tod80b] John Todd. John Hamilton Curtiss, 1909–1977. Annals of the History of Computing, 2(2):104–110, April/ June 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1980/pdf/a2104.pdf;   http:
//www.computer.org/annals/
an1980/a2104abs.htm.

Todd:1983:O

[Tod83] John Todd. Oberwolfach — 1945. In Beckenbach and Walter [BW83], pages 19–22. ISBN 3-7643-1539-3. LCCN QA295 .I57 1981. URL https://

link.springer.com/chapter/
10.1007/978-3-0348-6290-5_
2.

Todd:1984:ATT

[Tod84] John Todd. Applications of transformation theory: a legacy from Zolotarev (1847–1878). In S. P. Singh, J. W. H. Burry, and B. Watson, editors, Approximation theory and spline functions, Proceedings of the NATO Advanced Study Institute, St. John’s, Newfoundland 1983, number C 136 in NATO ASI Series, pages 207–245. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-6466-8.

Todd:1986:BPA

[Tod86] John Todd. The best polynomial approximation to (1 + x)^{−1} in [0, 1]. Adv. Math., Suppl. Stud., 10:313–320, 1986.

Todd:1987:PAH

[Tod87] J. Todd. The prehistory and ancient history of computation at the U.S. National Bureau of Standards. In Crane [Cra87], pages 47–51. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Todd:1988:LZ

[Tod88a] John Todd. A legacy from E. I. Zolotarev (1847–1878). The Mathematical Intelligencer, 10(2):50–53, 1988. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Todd:1988:MLM

[Tod88b] John Todd. The many limits of mixed means. II. the Carlson sequences. Numerische Mathematik, 54(1):1–18, October 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Todd:1990:PEH

[Tod90a] John Todd. The prehistory and early history of computation at the U.S. National Bureau of Standards. In Nash [Nas90], pages 251–268. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Todd:1990:WMP

[Tod90b] John Todd. The Weierstrass mean. I. the periods of ℘(z|e_{1}, e_{2}, e_{3}). Numerische Mathematik, 57(8):737–746, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Todd:1993:CQD

[Tod93] John Todd. Comments, queries, and debate: Alwin Walther (1898–1967). IEEE Annals of the History of Computing, 15(4):5–6, October/December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an1993/pdf/a4005.pdf.

Todd:1994:GHH

[Tod94a] John Todd. G. H. Hardy as an editor. The Mathematical Intelligencer, 16(2):32–37, 1994. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Todd:1994:RCL

[Tod94b] John Todd. Reminiscences of Cornelius Lanczos. In Brown et al. [BCEP94], pages lviii– lix. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Todd:2000:LC

[Tod00] John Todd. The lemniscate constants. In Berggren et al.

[BBB00], pages 412–417. ISBN 0-387-98946-3 (hardcover). LCCN QA484 .P5 2000. Reprint of [Tod75b].

Toole:1963:BRS

[Too63] Beverly M. Toole. Book review:

A Survey of Numerical Analysis by John Todd. Pi Mu Epsilon Journal, 3(8):422, Spring 1963. CODEN PMEJBR. ISSN 0031-952X. URL https://www.

jstor.org/stable/24338238.

Todd:1947:ACS

[TS47] John Todd and D. H. Sadler.

Admiralty Computing Service. Mathematical Tables and Other Aids to Computation, 2(19): 289–297, July 1947. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL https://www.
ams.org/journals/mcom/1947-
02-019/S0025-5718-47-99582-
3;   https://www.jstor.org/
stable/2002588.

Todd:1985:CPS

[TS85] John Todd and Jean E. Sammet. Comment on previous selfstudy answer and response. Annals of the History of Computing, 7(1):65, 68–69, January/ March 1985. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1985/pdf/a1065.pdf;   http:
//www.computer.org/annals/
an1985/a1065abs.htm.

Taussky:1940:CAN

[TT40a] Olga Taussky and John Todd.

A characterisation of algebraic numbers. Proceedings of the Royal Irish Academy. Section A. Mathematical and Physical Sciences, 46:1–8, 1940. ISSN 0035-8975. URL https://www.

                                        jstor.org/stable/20490743.

Taussky:1940:DQ

[TT40b] Olga Taussky and John Todd. Determinants of quaternions. Bulletin of the American Mathematical Society, 46(5):431–432, May 1940. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http:/

                                        /www.ams.org/journals/bull/
                                        1940-46-05/S0002-9904-1940-
                                        07220-9/S0002-9904-1940-07220-
                                        9.pdf.

Taussky:1940:MFP

[TT40c] Olga Taussky and John Todd. Matrices with finite period. Proceedings of the Edinburgh Mathematical Society (2), 6:128–134, 1940. CODEN PEMSA3. ISSN 0013-0915,1464-3839.

Taussky:1941:IG

[TT41a] Olga Taussky and J. Todd. Inversion in groups. Quarterly Journal of Mathematics. Oxford series., 12:65–67, 1941. CODEN QJMAAT. ISSN 0033-5606,1464-3847.

Taussky:1941:MFP

[TT41b] Olga Taussky and John Todd. Matrices of finite period. Proceedings of the Royal Irish Academy. Section A. Mathematical and Physical Sciences, 46: 113–121, 1941. ISSN 0035-8975.

URL https://www.jstor.org/
stable/20490752.

Taussky:1942:IPM

[TT42] Olga Taussky and John Todd. Infinite powers of matrices. Journal of the London Mathematical Society, 17(3):146–151, July 1942. CODEN JLMSAK. ISSN 0024-6107,1469-7750.

Taussky:1947:SAM

[TT47] Olga Taussky and John Todd.

Some aspects of modern algebra. Science Progress, 138:253–268, 1947. ISSN 0036-8504,2047-7163.

URL https://www.jstor.org/
stable/43412962.

Taussky:1948:CTGa

[TT48a] Olga Taussky and John Todd.

Covering theorems for groups. Bulletin of the American Mathematical Society, 54(3):276, March 1948. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1948:CTGb

[TT48b] Olga Taussky and John Todd.

Covering theorems for groups. Annales de la Societé Polonaise de Mathématique, 21:303– 305 (1949), 1948.

Taussky:1949:CTG

[TT49] Olga Taussky and John Todd.

Covering theorems for groups. Annales de la Société Polonaise de Mathématique, 21:303– 305, 1949.

Taussky:1951:CRP

[TT51] Olga Taussky and John Todd. The characteristic roots of products of matrices with rational integral elements. Bulletin of the American Mathematical Society, 57(2):123, March 1951. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1952:SEMa

[TT52a] Olga Taussky and John Todd. Systems of equations, matrices and determinants. I. Mathematics Magazine, 26(2):9–20, September/October 1952. CO-DEN MAMGA8. ISSN 0025-570X (print), 1930-0980 (elec-

tronic). URL https://www.
jstor.org/stable/3029837.

Taussky:1952:SEMb

[TT52b] Olga Taussky and John Todd. Systems of equations, matrices and determinants. II. Mathematics Magazine, 26(2):71–88, September/October 1952. CO-DEN MAMGA8. ISSN 0025-570X (print), 1930-0980 (elec-

tronic). URL https://www.
jstor.org/stable/3029698.

Taussky:1954:GTP

[TT54] Olga Taussky and John Todd. Generation and testing of pseudo-random numbers. Report 3370, U.S. National Bureau of Standards, Gaithersburg, MD, USA, June 22, 1954. ii + 16 pp. URL https:/

/nvlpubs.nist.gov/nistpubs/
Legacy/RPT/nbsreport3370.pdf.

Taussky:1956:CBT

[TT56a] Olga Taussky and John Todd.

Commuting bilinear transformations and matrices. Journal of the Washington Academy of Sciences, 46(12):373–375, December 1956. CODEN JWASA3. ISSN 0043-0439. URL https://www.

jstor.org/stable/24533908.

Taussky:1956:GTP

[TT56b] Olga Taussky and John Todd.

Generation and testing of pseudo-random numbers. In Symposium on Monte Carlo methods, University of Florida, 1954, pages 15–28. Wiley, New York, NY, USA, 1956.

Taussky:1957:ASE

[TT57a] Olga Taussky and John Todd.

Applications of systems of equations, matrices and determinants. In James and Beckenbach [JB57], chapter 24, pages 310–337. LCCN QA37 .J26.

URL https://www.questia.
com/read/86015503/the-tree-
of-mathematics.

Taussky:1957:SEM

[TT57b] Olga Taussky and John Todd.

Systems of equations, matrices and determinants. In James and Beckenbach [JB57], chapter 23, pages 305–318. LCCN

QA37 .J26. URL https://www.
questia.com/read/86015489/
the-tree-of-mathematics.

Taussky:1960:SDV

[TT60] Olga Taussky and John Todd.

Some discrete variable computations. In Proceedings of the Symposium on Applied Mathematics, volume 10, pages 201–209. American Mathematical Society, Providence, RI, USA, 1960.

Taussky:1991:ALM

[TT91] Olga Taussky and John Todd. Another look at a matrix of Mark Kac. Linear Algebra and its Applications, 150:341–360, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Taussky:2006:CTT

[TT06] Olga Taussky and John Todd. Cholesky, Toeplitz and the triangular factorization of symmetric matrices. Numerical Algorithms, 41(2):197–202, February 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=1017-1398&
volume=41&issue=2&spage=197.

Taussky:2012:ASE

[TT12a] Olga Taussky and John Todd. Applications of systems of equations, matrices and determinants. In James and Beckenbach [JB12], chapter 24, pages 310– 337. ISBN 1-258-28562-2. LCCN QA37 .J26.

Taussky:2012:SEM

[TT12b] Olga Taussky and John Todd. Systems of equations, matrices and determinants. In James and Beckenbach [JB12], chapter 23, pages 305–318. ISBN 1-258-28562-2. LCCN QA37 .J26.

Taussky-Todd:1957:ASE

[TTT57] Olga Taussky-Todd and John Todd. Applications of systems of equations, matrices, and determinants. Mathematics Magazine, ??(??), 1957. CODEN MAMGA8. ISSN 0025-570X. Special issue: “The Tree of Mathematics”.

Todd:1999:HIU

[TV99] John Todd and Richard S. Varga.

Hardy’s inequality and ultrametric matrices. Linear Algebra and its Applications, 302–303(1–3): 33–43, December 1, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/30/19/117/25/
28/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/
30/19/117/25/28/article.pdf;
 http://www.math.kent.edu/~
varga/pub/paper_227.pdf. Spe-

cial issue dedicated to Hans Schneider (Madison, WI, 1998).

Todd:1955:SLG

[TW55] John Todd and S. E. Warschawski.

On the solution of the Lichtenstein– Gershgorin integral equation in conformal mapping. II. Computational experiments. In Experiments in the computation of conformal maps, volume No. 42 of National Bureau of Standards Applied Mathematics Series, pages 31–44. United States Government Printing Office, Washington, DC, USA, 1955.

Williams-Hedges:1997:PCM

[WH97] Deborah Williams-Hedges. Pioneer of 20th-century mathematics John Todd dies. Caltech Media Relations Web site., June 25, 1997. URL http://

                                        mr.caltech.edu/media/Press_
                                        Releases/PR13007.html.

Yin:1994:CIF

[Yin94] Xiangrong Yin. A converse inequality of Fan, Taussky, and Todd. Journal of Mathematical Analysis and Applications, 182(3):654–657, March 15, 1994. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).